Actual costs for a company for the past 2 years were as follows:

Direct materials $9.00 per unit

Direct labour $8.00 per unit

Variable overhead $4.00 per unit

Variable selling and admin $3.00 per unit

Fixed overhead $80,000

Fixed selling and admin $315,000

During the year, 50,000 units were produced and 47,000 units were sold. There were 3000 units in beginning inventory.

a) What is the difference between absorption and variable net income? If absorption net income is $100,000 variable net income would be?

Variable net income=

Difference between absorption and variable net income=

b) Without preparing new income statements, what would be the difference in net income if sales were equal to production at 50,000 units and costs do not change from one year to the next?
Answer #1

a) Variable net income = 100000-(80000/50000*3000) = 95200

Difference in net income = (6000-3000)*1.6 = 4800

b) Difference in Net income = (3000-0)*1.6 = 4800